37.5% as a common fraction...how do you figure this out without a calculator??
Found 3 solutions by stanbon, onespal, gwmath:
Answer by stanbon(48545)

37.5% as a common fraction
Write as 37.5/100
Then multiply numerator and denominator by 10 to get:
375/1000
Cheers,
Stan H.

Answer by onespal(2)

I think the most simple process is to move the decimal to right...
one place to the right--- /10
two places to the right---- /100
3 -------------------------/1000
and so on..
then, just simplify the fraction.

Answer by gwmath(1)

37.5% as a common fraction without using a calculator:

Concept: There is a relationship between percentages, fractions, and decimals that help with conversions.
.25 = 25% = 1/4 ( think in terms of money: a quarter dollar is $.25, 25 cents expressed as a dollar = 25 cents/ 100 cents = 25% or simplifies to 1/4 as a common fraction)
.5 = 50% = 1/2
.75= 75% = 3/4
Solution:
37.5% = 37.5 parts out of a 100
= 37.5/100
= 37.5 x 10/ 100 x 10 ( Multiply the numerator and denominator by 10)
= 375/1000
= 75 / 200 ( simplify divide numerator and denominator by 5)
= 15 / 40
= 3 / 8